Kaspar Corporation makes a commercial-grade cooking griddle. The following information is available for Kaspar Corporation's anticipated annual volume of 25,600 units



Per UnitTotal
Direct materials$18
Direct labor$5
Variable manufacturing overhead$16
Fixed manufacturing overhead
$358,400
Variable selling and administrative expenses$3
Fixed selling and administrative expenses
$51,200

The company uses a 43% markup percentage on total cost.


(a)

 Compute the total cost per unit. Total cost per unit = $_______ 

Compute the target selling price.  Target selling price = $_______ 
Answer #1

a)Total cost per unit= direct material+direct labour+Var mfg overhead+var sG&A+ fixed mfg overhead+fixed sG&A
fixed mfg overhead per unit=fixed mfg overhead cost/no of units
=358400/25600=14
fixed SG&A per unit=Fixed seling and admin exp/no of units
=51200/25600=2
Total cost per unit=18+5+16+14+3+2=58

2)target selling prices=cost per unit+(markup*cost per unit)
=58+(43%*58)=82.94